Trial Concentration Level Oxygen Concentration Change Data (ppt) Elapsed Time (s) Enzyme Reaction Rate (ppt/s) 1...

Trial

Concentration

Level

Oxygen

Concentration

Change Data (ppt)

Elapsed

Time

(s)

Enzyme

Reaction Rate

(ppt/s)

1

2.5%

4.23

20.0

.21

2

5.0%

8.46

20.0

.42

3

10%

16.92

20.0

.85

                       

Observations and Questions

[1] Sample Calculation: Show your enzyme reaction rate calculation for Trial 1.

[2] Convert enzyme concentrations of 2.5%, 5.0%, and 10% to ppt units.
Answer #1

[1] As given in the question for trial 1

Oxygen concentration change data(ppt) = 4.23

Elapsed time(s) = 20

So to calculate enzyme reaction rate, we have to use the formula

Rate of reaction(ppt/s) = amount of concentration change/elapsed time

= 4.23/20

=0.21

[2] To convert the enzyme concentration of 2.5% , 5% and 10% we just have to divide the concentration levels by 100

For e.g,

(1) 2.5% = 2.5/100 = 0.025 ppt

(2) 5% = 5/100 = 0.05 ppt

(3) 10% = 10/100 = 0.1 ppt